Solve the following differential equations.

(i) dx/y²(x-y) =dy/x²(x-y) =dz/z(x²+y²)

(ii) (yz+z²)dx - xzdy + xydz =0

(iii) (z+z³)cosx dx - (z+z³)dy + (1-z²) (y-sinx)dz =0
1
Expert's answer
2019-09-02T09:11:56-0400

I)"\frac{dx}{y^2(x-y)}=\frac{dy}{x^2(x-y)}\Rightarrow x^2dx=y^2dy\Rightarrow x^3-y^3=C^3" . We have "\frac{dx}{y^2(x-y)}=\frac{dz}{z(x^2+y^2)}\Rightarrow \frac{(x^2+y^2)}{y^2(x-y)}dx=\frac{dz}{z}." Let "x=C\cosh^\frac{2}{3}t" and "y=C\sinh^\frac{2}{3}t" , then "dx=\frac{2C\sinh tdt}{3\cosh^\frac{1}{3}t}" and "\frac{(x^2+y^2)}{y^2(x-y)}dx=" "=\frac{C^2(\cosh^\frac{4}{3}t+\sinh^\frac{4}{3}t)}{C^2\sinh^\frac{4}{3}t\cdot C(\cosh^\frac{2}{3}t-\sinh^\frac{2}{3}t)}\cdot\frac{2C\sinh tdt}{3\cosh^\frac{1}{3}t}=\frac{2}{3}\frac{\cosh^\frac{4}{3}t+\sinh^\frac{4}{3}t}{\sinh^\frac{1}{3}t\cosh^\frac{1}{3}t(\cosh^\frac{2}{3}t-\sinh^\frac{2}{3}t)}dt" . Divide the numerator and denominator by "\cosh^\frac{4}{3}t" : "\frac{2}{3}\frac{\cosh^\frac{4}{3}t+\sinh^\frac{4}{3}t}{\sinh^\frac{1}{3}t\cosh^\frac{1}{3}t(\cosh^\frac{2}{3}t-\sinh^\frac{2}{3}t)}dt=\frac{2}{3}\frac{1+\tanh^\frac{4}{3}t}{\tanh^\frac{1}{3}t(1-\tanh^\frac{2}{3}t)}dt" Make change of the variable: "\tanh^\frac{2}{3}t=u" . So "t=artanh(u^\frac{3}{2})" and "dt=\frac{1}{1-u^3}\cdot\frac{3}{2}\sqrt{u}du" . Then"\frac{2}{3}\frac{1+\tanh^\frac{4}{3}t}{\tanh^\frac{1}{3}t(1-\tanh^\frac{2}{3}t)}dt=\frac{2}{3}\frac{1+u^2}{\sqrt{u}(1-u)}\cdot\frac{1}{1-u^3}\cdot\frac{3}{2}\sqrt{u}du=\frac{1+u^2}{(1-u)(1-u^3)}du" Write the partial fraction decomposition of "\frac{1+u^2}{(1-u)(1-u^3)}" : "\frac{1+u^2}{(1-u)(1-u^3)}=\frac{A}{1-u}+\frac{B}{(1-u)^2}+\frac{Du+E}{u^2+u+1}=" "=\frac{(A-Au+B)(u^2+u+1)+(Du+E)(1-u^2)}{(1-u)^2(u^2+u+1)}" "(A-Au+B)(u^2+u+1)+(Du+E)(1-u^2)=(-A+D)u^3+(B-2D+E)u^2+(B+D-2E)u+(A+B+E)" Equality of the polynoms "(-A+D)u^3+(B-2D+E)u^2+(B+D-2E)u+(A+B+E)" and "1+u^2" gives us system of equations: "\begin{cases}
-A+D=0\\
B-2D+E=1\\
B+D-2E=0\\
A+B+E=1
\end{cases}" , where "A=D=0, B=\frac{2}{3}" and "E=\frac{1}{3}" . So "\frac{1+u^2}{(1-u)(1-u^3)}=\frac{2}{3}\frac{1}{(1-u)^2}+\frac{1}{3}\frac{1}{u^2+u+1}" ."\int\bigl(\frac{2}{3}\frac{1}{(1-u)^2}+\frac{1}{3}\frac{1}{u^2+u+1}\bigr)du=\frac{2}{3}\int\frac{1}{(1-u)^2}du+\frac{1}{3}\int\frac{1}{u^2+u+1}du=\frac{2\sqrt{3}}{9}\arctan\bigl(\frac{2}{\sqrt{3}}u+\frac{1}{\sqrt{3}}\bigr)+\frac{2}{3(1-u)}+F" .Since "\frac{1+u^2}{(1-u)(1-u^3)}du=\frac{dz}{z}" , we have "\frac{2\sqrt{3}}{9}\arctan\bigl(\frac{2}{\sqrt{3}}u+\frac{1}{\sqrt{3}}\bigr)+\frac{2}{3(1-u)}+F = \ln |z|" or "e^Fe^{\frac{2\sqrt{3}}{9}\arctan\bigl(\frac{2}{\sqrt{3}}u+\frac{1}{\sqrt{3}}\bigr)+\frac{2}{3(1-u)}} = |z|" . That is "z=e^Fe^{\frac{2\sqrt{3}}{9}\arctan\bigl(\frac{2}{\sqrt{3}}u+\frac{1}{\sqrt{3}}\bigr)+\frac{2}{3(1-u)}}" or "z=-e^Fe^{\frac{2\sqrt{3}}{9}\arctan\bigl(\frac{2}{\sqrt{3}}u+\frac{1}{\sqrt{3}}\bigr)+\frac{2}{3(1-u)}}" . Union of these 2 cases is "z=Ge^{\frac{2\sqrt{3}}{9}\arctan\bigl(\frac{2}{\sqrt{3}}u+\frac{1}{\sqrt{3}}\bigr)+\frac{2}{3(1-u)}}" . Remembering that "u=\tanh^\frac{2}{3}t" , we can write "z=Ge^{\frac{2\sqrt{3}}{9}\arctan\bigl(\frac{2}{\sqrt{3}}\tanh^\frac{2}{3}t+\frac{1}{\sqrt{3}}\bigr)+\frac{2}{3(1-\tanh^\frac{2}{3}t)}}" Answer: "x=C\cosh^\frac{2}{3}t" , "y=C\sinh^\frac{2}{3}t" and "z=Ge^{\frac{2\sqrt{3}}{9}\arctan\bigl(\frac{2}{\sqrt{3}}\tanh^\frac{2}{3}t+\frac{1}{\sqrt{3}}\bigr)+\frac{2}{3(1-\tanh^\frac{2}{3}t)}}"

II)Dividing the equation by "xy^2" , we have "\frac{(yz+z^2)dx}{xy^2}+\frac{ydz-zdy}{y^2}=0" . Since "\frac{ydz-zdy}{y^2}=d\bigl(\frac{z}{y}\bigr)" , we can rewrite the equation: "\bigl(\frac{z}{y}+\bigl(\frac{z}{y}\bigr)^2\bigr)\frac{dx}{x}+d\bigl(\frac{z}{y}\bigr)=0" . After making change of the variables "\frac{z}{y}=t" , we have "(t+t^2)\frac{dx}{x}+dt=0" or "\frac{dx}{x}=-\frac{dt}{t+t^2}=\frac{dt}{t+1}-\frac{dt}{t}" . Solution of this equation is "\ln |x|+C=\ln |t+1|-\ln |t|=\ln\bigl |\frac{t+1}{t}\bigr|" . Rewrite this equaliy: "e^c|x|=\bigl|\frac{t+1}{t}\bigr|" . That is "e^cx=\frac{t+1}{t}" or "-e^cx=\frac{t+1}{t}" . Union of these 2 cases is "Dx=\frac{t+1}{t}" . Remembering that "t=\frac{z}{y}" , we have "Dx=\frac{y+z}{z}"

Answer: "Dx=\frac{y+z}{z}"

III)Rewrite the equation: "(z+z^3)(\cos xdx-dy)+(1-z^2)(y-\sin x)dz=0" . Note that "\cos xdx-dy=d(\sin x-y)" .

So after making change of the variables "\sin x-y=t" , we have "(z+z^3)dt-(1-z^2)tdz=0" or "\frac{dt}{t}=\frac{(1-z^2)dz}{z+z^3}"

"\frac{1-z^2}{z+z^3}=\frac{A}{z}+\frac{Bz+C}{z^2+1}=\frac{A(z^2+1)+z(Bz+C)}{z^3+z}=\frac{(A+B)z^2+Cz+A}{z^3+z}" , so "\begin{cases}
A+B=-1\\
C=0\\
A=1
\end{cases}" , where "A=1, B=-2, C=0"

We have "\frac{1-z^2}{z+z^3}=\frac{1}{z}-\frac{2z}{z^2+1}" . Since "\frac{dt}{t}=\frac{dz}{z}-\frac{2zdz}{z^2+1}" , we have "\ln |t| = \ln |z| - \ln |z^2+1| +C" or "|t|=e^C\bigl|\frac{z}{z^2+1}\bigr|" . That is "t=e^C\frac{z}{z^2+1}" or "t=-e^C\frac{z}{z^2+1}" . Union of these 2 cases is "t=D\frac{z}{z^2+1}" . Remembering that "t=\sin x-y" , we have "\sin x-y=D\frac{z}{z^2+1}"

Answer: "\sin x-y=D\frac{z}{z^2+1}"